OpenStudy (anonymous):

A garden has the shape of a right triangle with one leg 6 meters longer than the other. The hypotenuse is 6 meters less than twice the length of the shorter leg. Find the length of the shorter leg.

OpenStudy (anonymous):

let x be the shortest leg. One side is x, and the other is x + 6. The hypotenuse is 2x - 6. \[(2x-6)^{2} = x ^{2} + (x+6)^{2}\]

OpenStudy (anonymous):

Can you solve for x?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Too bad

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

4^2-24x+36 = 2x^2+12x+36 Can you solve it from there?

OpenStudy (anonymous):

all I did, by the way, was expand out everything that was in parentheses

OpenStudy (anonymous):

=2x^2+12x+36=4x^2-24x+36 0=2x^2-36x 2x^2-36=0 2(x^2-18)=0 x^2-18=0 x(x-18)=0 x=0 x-18=0 x=18 Length of the shorter leg is 18 meters Correct?